here's the sequence:
12, -4, 4/3, -4/9
what is the next two numbers??(or give me the pattern and i'll figure them out)
thanks

Answer by checkley71(8403)

12(-1/3)=-4(-1/3)=4/3(-1/3)=-4/9 ETC.